### v3

@@ Abstract
Many studies have shown how ergonomically designed furniture improves productivity and well-being. As computers have become a part of students’ academic lives, they will grow further in the future. We propose anthropometric-based furniture dimensions suitable for university students to improve computer laboratory ergonomics. We collected data from 380 participants and analyzed 11 anthropometric measurements, correlating them to 11 furniture dimensions. Two types of furniture were found and studied in different university computer laboratories: (1) a non-adjustable chair with a non-adjustable table and (2) an adjustable chair with a non-adjustable table. The mismatch calculation showed a significant difference between existing furniture dimensions and anthropometric measurements, indicating that 7 of 11 existing furniture dimensions need improvements. The one-way ANOVA test with a significance level of 5% also showed a significant difference between anthropometric data and existing furniture dimensions. All 11 dimensions were determined to match the student’s anthropometric data. The proposed dimensions were found to be more compatible and showed reduced mismatch percentages for 9 furniture dimensions and nearly zero mismatches for SW, BH, and UTH for both males and females compared to existing furniture dimensions. The proposed dimensions of the furniture set with adjustable seat height showed slightly improved results for SH and STC, which showed zero mismatches compared to the non-adjustable furniture set. The TW and TD dimensions were suggested according to Barnes and Squires’ ergonomic work envelope model, considering hand reach. The position for the keyboard and mouse was also suggested according to the work envelope. The monitor position and viewing angle were proposed according to OSHA guidelines. This study suggests that the proposed dimensions can improve comfort levels, reducing the risk of musculoskeletal disorders among students. Further studies on the implementation and long-term effects of these proposed dimensions in real-world computer laboratory settings are recommended.
@@ 1 Introduction
The use of computers in professional environments is experiencing an upward trend. Moreover, the use of home computers for both occupational and educational purposes has seen a notable surge over the preceding decades (Py Szeto et al.,, 2014). Poor posture linked to furniture is a recognized factor that can increase the likelihood of developing musculoskeletal disorders (MSDs) (Magnusson and Pope,, 1998). Therefore, human body measurements are the most critical factor in designing furniture that can effectively adapt to posture, thereby enhancing comfort (Kopardekar,, 1996). Anthropometry plays a vital role in designing work environments to reduce the occurrence of MSDs among employers and employees, ultimately leading to increased productivity (Afzan et al.,, 2012). A study was conducted on the postural angles of primary body segments, including the neck, elbow, and wrist, among six sets of parents and children while using a computer at home (Williams and Jacobs,, 2002). Ergonomics seeks to create work environments that allow individuals to work easily, efficiently, and safely (Savanur et al.,, 2007). Creating equipment, merchandise, furniture, and other devices that match the human body’s needs can enhance the comfort, physical health, well-being, and performance of individuals (Parvez et al.,, 2018).
@@
The motivation for this study stems from the growing recognition of the importance of ergonomic design in enhancing human productivity and well-being. In recent years, numerous studies have demonstrated the positive impact of ergonomically designed furniture on comfort, health, and performance, particularly in settings where individuals spend prolonged periods of time seated at desks, such as offices and educational institutions. As computers and digital technology continue to permeate various aspects of academic and professional life, ensuring optimal ergonomic conditions in computer labs has become increasingly crucial. While ergonomic guidelines and standards exist for general office environments, there is a notable gap in tailored guidelines for computer labs, particularly for university students.
The prevalence of computers in educational settings, especially universities, highlights the need to prioritize students’ physical comfort and health by providing ergonomically sound furniture. However, the lack of specific guidelines for computer laboratories presents a challenge in achieving this objective. This study addresses this gap by proposing anthropometric-based furniture dimensions that are specifically designed for university students. By analyzing data collected from 380 participants and correlating 11 anthropometric measurements to 11 furniture dimensions, this research aimed to optimize furniture design in university computer labs to better accommodate students’ physical characteristics.
@@
We aimed to identify potential mismatches between students’ anthropometric measurements and computer lab furniture dimensions to establish better dimensions for the design parameters of computer lab furniture for university students, considering relevant anthropometric measurements of students sitting to minimize potential mismatches and prolonged health issues. This research aims to assess and compare the potential mismatch between two existing sets of computer tables and chairs and propose standardized design criteria and dimensions based on anthropometry and safety guidelines for students. This will involve analyzing 11 anthropometric data points along 11 furniture dimensions using ANOVA for decision-making. It is anticipated that aligning furniture more closely with the human body will result in a notable reduction in MSDs and an associated increase in productivity. Both non-adjustable and adjustable furniture dimensions must be adjusted according to the anthropometric data of the undergraduate students.
This study focuses on optimizing university furniture design using personalized anthropometric measurements tailored to university student populations. By analyzing 11 key anthropometric measurements and comparing them with existing furniture, I provide a comprehensive understanding of students’ ergonomic needs. ANOVA statistical analysis adds robust validation to my findings, ensuring that any proposed ergonomic design is grounded in solid statistical evidence. My research proposes innovative and ergonomically designed furniture solutions based on current student body measurement data, potentially improving student comfort, productivity, and overall well-being. These findings have significant implications for educational institutions and furniture manufacturers, highlighting the importance of customizing furniture designs to meet users’ specific needs. My research contributes to the ergonomics and design fields by addressing the literature gap regarding the ergonomic needs of university students. The proposed ergonomic designs have practical applications in real-world settings, such as university classrooms, computer labs, and study areas, where they can enhance the learning environment and promote student health and well-being. This research also opens the door for further exploration of the ergonomic needs of different user groups and the development of more tailored furniture solutions for various settings.
The contributions of the proposed work are given below:
  1. 1.
  2. 2.
    By identifying the mismatch between existing furniture dimensions and anthropometric measurements, this research offers practical insights for designing ergonomic furniture in university environments, improving comfort, health, and productivity among students and staff.
  3. 3.
  4. 4.
    We collected and analyzed anthropometric data from university students, providing valuable insights into the body measurements of the target population.
  5. 5.
    We conducted a correlation analysis among 11 anthropometric measurements, shedding light on the interrelationships between different body dimensions.
  6. 6.
    By comparing the mismatch between two different types of existing furniture and anthropometric data, the study offers a nuanced understanding of how different furniture designs affect users’ comfort and well-being.
  7. 7.
    The study employed ANOVA for better decision-making, providing a robust statistical framework for assessing the significance of differences among existing furniture types and anthropometric measurements.
  8. 8.
    Drawing from the current students’ body measurement data, this study proposes ergonomically designed furniture that aligns with the specific needs and dimensions of the target population, offering practical solutions for enhancing comfort and productivity in university settings.
The paper is structured as follows: The “Introduction” section overviews the research problem and outlines the significance of addressing ergonomic challenges in university environments. In the “Literature Review” section, we review existing studies on ergonomics and anthropometric measurements in university settings, highlighting the gaps that our research aims to fill. The “Methodology” section details our approach to collecting anthropometric data and designing furniture based on these measurements. In the “Results and Discussion” section, we present the findings of our anthropometric survey, evaluate the proposed ergonomic furniture design, interpret and discuss the results. The “Conclusions” section summarizes the main findings and contributions of the study and proposes future research directions.
@@ 2 Literature Review
Self et al., (2019) conducted a study that examined the potential factors that promote or hinder interdisciplinary education in design, specifically from the viewpoint of student learning. The researchers conducted a series of detailed interviews with students from two interdisciplinary design schools, one located in the Republic of Korea and the other in the UK. This study discusses the implications of adopting interdisciplinary approaches in design education and proposes areas for further research in this domain (Self et al.,, 2019). Laeser et al., (1998) recognized the necessity for furniture designed for 6th and 8th grade children. In their study, students from these grades performed keyboarding and mousing tasks at two distinct computer workstations. One workstation was a conventional desktop setup, whereas the other featured an adjustable keyboard. The study found that the overall posture scores of the students improved when using the adjustable workstation, as indicated by the Rapid Upper Limb Assessment. Thariq, (2010) found that current side-mounted desktop chair designs are not suited to fulfilling the postural and comfort requirements of university students in their learning environment. In response, the researchers conducted an anthropometric survey of university students in Sri Lanka. The survey aimed to create fixed-type and side-mounted desktop chairs and evaluate their comfort levels based on design dimensions. This research emphasizes the significance of chairs that support the development of a healthy and comfortable posture among students, which can contribute to creating a healthy workforce in various industries. Taifa and Desai, (2017) recognized the issue of poorly designed furniture for engineering students in India. The authors employed a health survey, specifically an ergonomic assessment, to gather and analyze the data in their study. Based on their findings, the authors proposed the design of adjustable classroom furniture as a solution to reduce MSDs.
Hoque et al., (2014) found that classroom furniture does not consider ergonomics. The authors conducted an anthropometric survey of 500 Bangladeshi university students, followed by a mismatch analysis. Based on their findings, the authors recommend the use of ergonomically fit furniture. Joshi et al., (2015) addressed the significant rise in the design of computer workstations and associated health and safety issues for users. Their research aimed to evaluate students’ knowledge at the G.B.P. University of Agriculture and Technology concerning the use of ergonomically appropriate computer workstations. The authors conducted a postural assessment using the participative ergonomic technique of Rapid Upper Limb Assessment (RULA). The study revealed that most computer users lacked adequate information regarding the positioning of computer workstations, and a substantial number of computer operators reported issues with their upper extremities. Mowatt et al., (2018) aimed to determine the prevalence of computer vision syndrome (CVS) and ergonomic practices among students in the Faculty of Medical Sciences at The University of the West Indies, Jamaica. A cross-sectional study was conducted for research purposes using a self-administered questionnaire. The study found that ocular symptoms and neck pain were less likely when the device was held just below eye level. The study highlighted a high prevalence of CVS symptoms among university students, suggesting that adopting improved ergonomic practices could potentially reduce issues such as neck pain, eye strain, and burning. Kahya, (2019) investigated the congruence between the dimensions of school furniture and anthropometric measurements of students. They assessed nine anthropometric dimensions of 225 students from nine departments. The results indicated a significant mismatch, with a seat height discrepancy of 44.45% discrepancy, seat depth of 100% discrepancy, and desk height of 21.28% discrepancy. This study lays the groundwork for and encourages further exploration of classroom ergonomics in university settings. Abd Rahman et al., (2018) developed an anthropometric database tailored to Malaysian operators, focusing on sitting and standing dimensions. The database was constructed from measurements of 146 male and 168 female participants aged 18 - 45 years. Thirty-six anthropometric dimensions were chosen to compare their common availability across the four countries. This knowledge of population variations is crucial for designing workstations and facilities that accommodate the needs of the industrial environment.
@@
Zayed et al., (2021) sought to understand the prevalence and factors associated with Digital Eye Strain (DES) among IT professionals at Tanta University, Egypt. They employed an interview questionnaire to collect and analyze data, revealing significant predictors of DES, including sex, age, wearing eyeglasses, prolonged computer use, an unadjusted ergonomic workstation, and a dry environment. The study concluded that DES can be mitigated by increasing knowledge and awareness through eye health education, regular training in ergonomic workstations, and adjustments to create a comfortable workplace environment. Parvez et al., (2018) identified that poorly fitting furniture could contribute to various MSDs and discomforts. To delve into this issue, they surveyed ten primary schools in Bangladesh, revealing a notable disparity between furniture dimensions and student body measurements. Moreover, this paper proposes furniture dimensions that could notably decrease the mismatch percentage among students from 90% to 10%. Shohel Parvez et al., (2022) acknowledged that inadequately designed furniture could potentially exacerbate MSDs. They conducted a survey of 400 university students (250 males and 150 females) who volunteered for the study. The Standard Nordic Musculoskeletal Questionnaire was used to evaluate MSDs. The authors recommended that modifying the design of academic furniture could help alleviate or prevent MSD.
Dessie et al., (2018) observed that prolonged computer use could increase the risk of CVS in users. They conducted a cross-sectional study to gauge the prevalence of CVS and its related factors among government employees in Debre Tabor town, encompassing 607 participants. The research revealed that computer users with a comprehensive understanding of computer use safety measures and associated adverse effects were less susceptible to CVS. Boadi-Kusi et al., (2020) emphasized the importance of addressing poor ergonomic office setups that can lead to university administrative staff in Ghana spending prolonged periods behind computer screens, thus increasing the risk of developing CVS. To address this issue, a cross-sectional survey was conducted among 200 administrative staff members at the University of Cape Coast. The results indicated that raising awareness about CVS and adhering to recommended ergonomic practices are crucial in reducing the prevalence of CVS and, consequently, enhancing work satisfaction and productivity. Moradi et al., (2017) conducted an ergonomic risk assessment of auto mechanics. This study found that the back and waist had the highest prevalence of work-related MSDs. Fazi et al., (2019) utilized the Rapid Upper Limb Assessment (RULA) method to assess risks in an automotive manufacturing company. These findings demonstrated an exceptionally high level of risk, highlighting the urgent need for immediate changes. Kibria, (2019) developed a self-report Nordic ergonomic assessment questionnaire to identify discrepancies between furniture dimensions and anthropometric measurements. The proposed ergonomically designed computer workstation, informed by anthropometric measurements and guidelines, could potentially alleviate MSDs among university teachers. This study could significantly influence ergonomic furniture design in universities and other organizations, addressing computer-related ergonomic challenges.
@@
The current literature on ergonomics and its impact on user health and performance has identified several research gaps that warrant further investigation. First, comprehensive research on interdisciplinary education in design and its impact on student learning outcomes is lacking. Second, few studies have been conducted on the impact of ergonomic furniture design on children’s posture and health. Third, there is a shortage of research on the ergonomic needs of university students, particularly in different countries and cultural contexts. Fourth, studies on the prevalence and impact of CVS among university students and professionals in various occupational settings are scarce. Fifth, there is a lack of research on the effectiveness of ergonomic interventions in reducing the risk of MSDs and other ergonomic-related issues in occupational settings. Sixth, limited research has been conducted on the impact of ergonomic interventions on improving work satisfaction, productivity, and overall well-being in various occupational settings. Seventh, there is a shortage of research on the long-term effects of ergonomic interventions on employees’ health, performance, and retention. Finally, there is a lack of studies on the cost-effectiveness of ergonomic interventions and their potential return on investment in different occupational settings. Addressing these research gaps will contribute to developing more effective ergonomic interventions and overall well-being in various occupational settings. However, there has been a limited emphasis on the design of furniture and computer labs for undergraduate education in universities, and there is a lack of established standards for computer lab furniture in Bangladeshi universities. Bangladeshi universities typically use desktops, chairs, and tables as computer lab furniture. Prolonged use of these improperly designed furniture pieces could exacerbate MSDs in young adults. Hence, ergonomically fitted furniture is crucial. Failure to address these issues in a timely manner could result in undetected severe health problems among the students.
3 Methodology
The main goal of this study is to assess the current state of computer laboratories in Bangladesh. This will be accomplished by gathering anthropometric data from the students. The objective was to establish standardized criteria and measurements for the design of computer tables in Bangladesh. The sequential steps of this study are outlined in Figure 1.
Figure 1: Flowchart illustrating the step-by-step methodology followed in this research.

### v4

@@ Abstract
Many studies have shown that ergonomically designed furniture improves productivity and well-being. As computers have become a part of students’ academic lives, they will continue to grow in the future. We propose anthropometric-based furniture dimensions that are suitable for university students to improve computer laboratory ergonomics. We collected data from 380 participants and analyzed 11 anthropometric measurements, correlating them with 11 furniture dimensions. Two types of furniture were found and studied in different university computer laboratories: (1) a non-adjustable chair with a non-adjustable table and (2) an adjustable chair with a non-adjustable table. The mismatch calculation showed a significant difference between existing furniture dimensions and anthropometric measurements, indicating that 7 of the 11 existing furniture dimensions need improvement. The one-way ANOVA test with a significance level of 5% also showed a significant difference between the anthropometric data and existing furniture dimensions. All 11 dimensions were determined to match students’ anthropometric data. The proposed dimensions were found to be more compatible and showed reduced mismatch percentages for nine furniture dimensions and nearly zero mismatches for seat width, backrest height, and under the hood for both males and females compared to the existing furniture dimensions. The proposed dimensions of the furniture set with adjustable seat height showed slightly improved match results for seat height and seat-to-table clearance, which showed zero mismatches compared with the non-adjustable furniture set. The table width and table depth dimensions were suggested according to Barnes and Squires’ ergonomic work envelope model, considering hand reach. The positions of the keyboard and mouse are also suggested according to the work envelope. The monitor position and viewing angle were proposed according to OSHA guidelines. This study suggests that the proposed dimensions can improve comfort levels, reducing the risk of musculoskeletal disorders among students. Further studies on the implementation and long-term effects of the proposed dimensions in real-world computer laboratory settings are recommended.
@@
[…]
[inst3]organization=Advanced Machine Intelligence Research (AMIR) Lab, city=Dhaka, postcode=1229, country=Bangladesh
@@ 1 Introduction
Ergonomics plays a crucial role in ensuring safety, health, and performance in various settings, including educational institutions, where efforts are made to optimize work and study environments for both teachers and students to enhance productivity and minimize the risk of musculoskeletal discomfort [1]. Integrating ergonomic principles into education can improve the quality and increase productivity. For instance, research has shown that sitting posture and positioning significantly influence typing and handwriting performance [2]. Designing ergonomic solutions often relies on tools such as anthropometry, which involves measuring the size, shape, and capabilities of the human body for task performance [3, 4].
Several studies have proposed methods and guidelines for designing ergonomic classroom furniture tailored to students of different ages, providing relevant anthropometric dimensions and data for such designs [5, 6]. Among the various models for ergonomic furniture design, the ’design for adjustable range’ model is widely recommended, as it considers a range of anthropometric measures of end-users, typically between the 5th and 95th percentiles of the user population [7].
@@
The motivation for this study stems from the growing recognition of the importance of ergonomic design for enhancing human productivity and well-being. Numerous studies have demonstrated the positive impact of ergonomically designed furniture on comfort, health, and performance, especially in settings where individuals spend prolonged periods of time seated at desks, such as offices and educational institutions. As computers and digital technology continue to permeate various aspects of academic and professional life, ensuring optimal ergonomic conditions in computer laboratories has become increasingly crucial. However, while ergonomic guidelines and standards exist for general office environments, tailored guidelines for computer laboratories, particularly for university students, are lacking.
This study addresses this gap by proposing anthropometric-based furniture dimensions specifically designed for university students. This research aims to optimize furniture design in university computer labs to better accommodate students’ physical characteristics.
@@
The main goal of this study was to assess the current status of computer laboratories in Bangladesh by gathering anthropometric data from students. We sought to identify potential mismatches between anthropometric measurements and computer lab furniture dimensions to establish better design parameters for university students. By considering relevant anthropometric measurements of students sitting, our aim was to minimize potential mismatches and prolonged health issues. Additionally, this study aims to assess and compare the potential mismatch between two existing computer tables and chair sets. We proposed standardized design criteria and dimensions based on anthropometry and student safety guidelines.
Implementing anthropometric-based furniture dimensions tailored for university students can significantly enhance the ergonomic design of computer laboratories. By reducing potential mismatches between students’ physical measurements and furniture dimensions, this research contributes to the creation of safer, more comfortable, and more productive learning environments. These optimized furniture dimensions have the potential to minimize the risk of musculoskeletal disorders (MSDs) and discomfort among students, thereby promoting their overall well-being and academic performance. Furthermore, by providing tailored guidelines for computer laboratory furniture design, this study offers practical insights that can inform future research and guide the development of ergonomic standards in educational institutions.
@@ 2 Literature Review
Hitka and Gejdoš analyzed a sample of Slovak university students from 2000-2006 and 2018-2023 for 25 anthropometric measures relevant to workspace design. They showed a decrease in Slovakia’s adult population trend, with significant changes observed only in body weight among young adults [8]. Kang et al. investigated changes in body dimensions over time in Western (US) and Eastern (Korean) populations, focusing on the dimensions pertinent to automobile driver seat design. Data from the ANSUR and Size Korea datasets at two time points showed a consistent increase in body dimensions for both sexes in both populations [9]. Rababah and Etier showed that anthropometric measures in children are vital for designing child-centric furniture, tools, and toys. They analyzed a dataset of 354 children from Jordan across six age groups (six months to nine years), including 23 anthropometric measurements [10]. Goleij et al. (2024) reviewed anthropometric changes worldwide over 30 years and analyzed 132 articles. They found significant increases in various body dimensions, emphasizing the need for regular updates and adaptations in work environments and equipment based on new anthropometric data [11]. Hitka et al. (2022) determined the ideal dimensions of wooden chairs for the adult bariatric subpopulation of Slovakia. By tracking long-term anthropometric changes, they were able to identify the body height of 95% of adult males [12].
@@
Laeser et al. recognized the necessity for furniture designed for 6th and 8th-grade children. In their study, students from these grades performed keyboarding and mousing tasks at two computer workstations. One workstation was a conventional desktop setup, and the other featured an adjustable keyboard. The study found that the overall posture scores of the students improved when using the adjustable workstation, as indicated by the Rapid Upper Limb Assessment [13]. Thariq found that the current side-mounted desktop chair designs are not suited to fulfilling the postural and comfort requirements of university students in their learning environment. They conducted an anthropometric survey of university students in Sri Lanka. The survey aimed to create fixed-type and side-mounted desktop chairs and evaluate their comfort levels based on the design dimensions [14]. Taifa and Desai recognized the issue of poorly designed furniture for engineering students in India. The authors employed a health survey, specifically an ergonomic assessment, to gather and analyze the data in their study. Based on their findings, the authors proposed the design of adjustable classroom furniture as a solution to reduce MSDs [15]. Hoque et al. found that classroom furniture does not consider ergonomics. The authors conducted an anthropometric survey of 500 Bangladeshi university students, followed by a mismatch analysis. Based on their findings, the authors recommended the use of ergonomically fit furniture [16]. Joshi et al. addressed the significant rise in the design of computer workstations and the associated health and safety issues for users. They conducted postural assessment using the participative ergonomic technique of Rapid Upper Limb Assessment (RULA). The study revealed that most computer users lacked adequate information regarding the positioning of computer workstations, and a substantial number of computer operators reported issues with their upper extremities [17].
Kahya investigated the congruence between school furniture dimensions and students’ anthropometric measurements. They assessed nine anthropometric dimensions of 225 students from nine departments. The results indicated a significant mismatch, with a seat height discrepancy of 44.45% discrepancy, seat depth of 100% discrepancy, and desk height of 21.28% discrepancy. This study lays the groundwork for further exploration of classroom ergonomics in university settings [18]. Abd Rahman et al. developed an anthropometric database tailored to Malaysian operators, focusing on sitting and standing dimensions. The database was constructed from the measurements of 146 male and 168 female participants aged 18-45 years. Thirty-six anthropometric dimensions were chosen to compare their availability across the four countries. This knowledge of population variations is crucial for designing workstations and facilities that accommodate the needs of the industrial environment [19]. Langová et al. presented a method for designing chair dimensions based solely on the height range of intended users using seven equations describing the relationship between key design dimensions and human height or height range [20]. Sydor and Hitka linked chair dimensions to anthropometric measurements, simplifying access to data by establishing average body proportions for adults. Through seven equations, they related chair design dimensions to human height, thereby facilitating design considerations [21].
@@
Parvez et al. identified that poorly fitting furniture could contribute to various MSDs and discomfort. To delve into this issue, they surveyed ten primary schools in Bangladesh, revealing a notable disparity between furniture dimensions and student body measurements. Moreover, this study proposes furniture dimensions that could notably decrease the mismatch percentage among students from 90% to 10% [22]. Shohel Parvez et al. acknowledged that inadequately designed furniture could potentially exacerbate MSDs. They conducted a survey of 400 university students (250 males and 150 females) who volunteered for the study. The Standard Nordic Musculoskeletal Questionnaire was used to evaluate MSDs. The authors recommended that modifying the design of academic furniture could help alleviate or prevent MSD [23]. Moradi et al. conducted an ergonomic risk assessment of auto mechanics. This study found that back and waist circumference had the highest prevalence of work-related MSDs [24]. Kibria and Rafiquzzaman developed a self-reported Nordic ergonomic assessment questionnaire to identify discrepancies between furniture dimensions and anthropometric measurements. The proposed ergonomically designed computer workstation, informed by anthropometric measurements and guidelines, can potentially alleviate MSDs among university teachers. This study could significantly influence ergonomic furniture design in universities and other organizations by addressing computer-related ergonomic challenges [25].
As computer use grows, the risks of physical strains and eye issues increase. To combat this, various international standards like the International Organization for Standardization [26] and guidelines from organizations such as the Occupational Safety and Health Administration [27] offer recommendations for ergonomic workstation design. These focus on factors such as the chair and desk setup, monitor positioning, and lighting to reduce strain and promote better health during computer use. The International Organization for Standardization standard specifies basic dimensions for furniture used in office work, including tables and chairs, to promote user comfort and safety [28]. Additionally, The European Committee for Standardization (2020) provided guidelines for the ergonomic design of workstations, including layout and postural requirements for tasks involving visual display terminals (VDTs). This standard offers valuable insights into the workstation layout and postural requirements, which are crucial considerations in computer laboratory furniture design [26]. Additionally, the European Committee for Standardization specifies the functional dimensions of the chairs and tables used in educational institutions. This European standard outlines ergonomic furniture requirements to promote students’ health and well-being [29]. European Committee for Standardization (2012) provides additional safety and ergonomic requirements for educational furniture, complementing [30, 31]. Similarly, the European Committee for Standardization (2000) sets safety and ergonomic requirements for office chairs, ensuring user comfort and reducing the risk of musculoskeletal disorders [32]. Incorporating the recommendations outlined in these standards ensures that computer laboratory furniture is ergonomically optimized for university students, minimizing the risk of MSDs and promoting student comfort and productivity.
@@
Current literature on ergonomics and its impact on user health and performance has identified several research gaps that warrant further investigation. First, comprehensive research on interdisciplinary education in design and its impact on student learning outcomes is lacking. Second, few studies have been conducted on the impact of ergonomic furniture design on children’s posture and health. Third, there is a shortage of research on university students’ ergonomic needs, particularly in different countries and cultural contexts. Fourth, research on the effectiveness of ergonomic interventions in reducing the risk of MSDs and other ergonomic-related issues in educational settings is lacking. Fifth, limited research has been conducted on the impact of ergonomic interventions on work satisfaction, productivity, and overall well-being in various educational settings. Sixth, there is a shortage of research on the long-term effects of ergonomic interventions on students’ health, performance, and retention. Finally, studies on the cost-effectiveness of ergonomic interventions and their potential return on investment in different occupational settings are lacking. Addressing these research gaps will contribute to the development of more effective ergonomic interventions and overall well-being in various occupational settings. However, there has been a limited emphasis on the design of furniture and computer labs for undergraduate education in universities, and there is a lack of established standards for computer lab furniture in Bangladeshi universities. Bangladeshi universities typically use desktops, chairs, and tables as computer laboratory furniture. The prolonged use of these improperly designed furniture pieces could exacerbate MSDs in young adults. Hence, ergonomically fitted furniture is crucial. Failure to address these issues promptly could result in undetected severe health problems among the students.
3 Materials and Methods